At a school fundraiser, Annie purchases some t-shirts as prizes. The medium sized shirts were $4 per shirt and the large were $5 per shirt. If she ordered a total of 70 shirts with a total cost of $320, how many t – shirts were size medium?


1
Expert's answer
2021-05-21T11:07:32-0400

"m+l=70;\;4m+5l=320."

"l=70-m."

"4m+5(70-m)=320."

"4m-5m=320-350."

"m=30."

30 t – shirts were size medium.